OpenStudy (meilendurcer):

6 yd, 1 ft = __ ft A. 7 B. 12 C. 19 D. 36

OpenStudy (goformit100):

. 19

OpenStudy (meilendurcer):

thank you

OpenStudy (goformit100):

Any more question?

OpenStudy (meilendurcer):

6 1/2 lb = __ oz A. 104 B. 108 C. 112 D. 136

OpenStudy (goformit100):

B. 108

OpenStudy (meilendurcer):

thanx

OpenStudy (goformit100):

Any more?

hero (hero):

6.5 lb = x oz replace lb with 16 oz 6.5 * 16 oz = x oz Divide both sides by oz 6.5*16 = x 104 @meilendurcer

OpenStudy (meilendurcer):

@Hero thanks! i wish i saw that answer yesterday before i put it in